In order to finish the order on time the factory was supposed to produce 70 items daily. After hiring three more people they were able to produce 10 more items daily, and completed the order five days ahead of the schedule. How many days was the order supposed to take originally?
Answer by ikleyn(52767)

.
70*x = (70+10)*(x-5)   <<<---===  It is what the condition says.


70x = 80*(x-5)


70x = 80x - 400


400 = 80x - 70x


10x = 400  ===========>>>   x = 400/10 = 40.


ANSWER.  40 days.

---------------

The info about 3 new hired workers is   a)  Excessive;   b)  Unnecessary;   c)  it is not used in the solution;
and   d)  it is included to confuse the reader and to distract his attention.